# Beyond Semantics: Rediscovering Spatial Awareness in Vision-Language Models

### v1

@@ 5.1 A 2D Synthetic Spatial Benchmark (2DS)
Standard image and text benchmarks often unintentionally provide semantic or contextual shortcuts, enabling models to leverage object co-occurrence or linguistic biases without genuinely understanding spatial relationships. We particularly hope to respond to the potential inadequacies in current benchmarks from our spatial compression study (Section 2). To isolate and rigorously evaluate spatial reasoning, we limit confounding factors such as real-world object frequencies and linguistic priors. Our benchmark specifically targets 2D spatial relationships directly influenced by positional embeddings.
Example Queries including:
  • “What color is at the bottom of the image?” (Color, absolute position)
  • “Is the circle below the square?” (Shape, relative position)
5.1.1 Dataset Construction Details
Figure 4 provides representative examples, emphasizing the simplicity of semantic attributes and the centrality of spatial relationships.
Figure 4: Illustrative scenes from our 2DS Dataset. The left two images show two object arrangements, while the right two images illustrate more complex three and four object arrangements.
5.2 Interpretability-Informed Model Adjustments
Our interpretability analyses identified embedding norm suppression and insufficient usage of local geometric information as critical limitations to effective spatial reasoning in VLMs. Guided by these insights, we propose two straightforward and interpretable adjustments: (1) Vision Token Normalization and (2) Utilizing Intermediate-Layer Features.
Vision Token Normalization Our prior analyses revealed that vision embedding magnitudes significantly exceed those of positional embeddings, effectively suppressing positional signals during attention computations. To address this imbalance:
  • Norm Matching: We measured the distribution of typical text embedding norms (mean $\approx 0.83$, max $\approx 1.22$).
  • RMS Normalization: We applied RMS normalization [39] that is used in LLM [6] to vision encoder outputs to match them closely with textual embedding norms.
This normalization reduces the dominance of semantic signals, allowing positional cues to emerge more clearly.
Intermediate-Layer Features (Local Geometry) Current VLM architectures typically use highly aggregated, semantically-rich final-layer vision features from CLIP, which excel in semantic alignment but may lack fine-grained spatial details [36]. Recent studies suggest that intermediate layers of CLIP encoders retain richer local spatial information beneficial for VLMs [18]. Thus, we investigate whether directly incorporating earlier spatial features improves spatial reasoning.
Specifically, we extracted features from intermediate vision encoder layers (12, 16, 20, and 24), inherently containing more detailed spatial information reminiscent of the dorsal-stream processing pathway. These mid-layer features explicitly encode fine-grained local spatial information, potentially crucial for accurately answering purely spatial queries in our synthetic dataset.
However, an interpretability-informed ablation study (Section 7.2) reveals a semantic-spatial trade-off: while intermediate features significantly boost purely spatial task performance, they might also reduce the LLM’s incentive to develop explicit positional sensitivity within its attention mechanism.
6 Experiments
Figure 5: Accuracy comparison across varying numbers of objects. Our interpretability-informed adjustments yield consistent improvements, especially as spatial complexity increases.
We now evaluate the effectiveness of our interpretability-informed interventions, namely vision embedding normalization and intermediate-layer feature extraction, in enhancing spatial reasoning in VLMs. Our experiments cover results on both a controlled synthetic spatial benchmark and several standard vision-language benchmarks.
6.1 Experimental Setup
We use LLaVA 1.5 (7B) [23] as our baseline model, which is chosen for its open-source availability and transparent training details. We replicate LLaVA 1.5’s exact two-stage training process and dataset usage for fair comparison. Based on the interpretability insights discussed earlier, we evaluate two proposed variants:
  • + Normalize: Vision embeddings normalized according to the procedure detailed in Section 5.2.
  • + Normalize + Multilayer: Normalized vision embeddings augmented with intermediate-layer features to retain fine-grained spatial information ( Section 5.2).
6.2 Results on Synthetic Dataset
Table 2 summarizes accuracy across distinct spatial query categories, and Figure 6 illustrates performance across varying numbers of objects, clearly highlighting the effectiveness of our interventions.
Figure 6: Accuracy comparison across different categories of 2DS dataset. We see consistent improvements across categories.
@@ 6.2 Results on Synthetic Dataset
Table 2: Spatial reasoning accuracy (%) across 2DS categories. $\uparrow$ indicates higher is better. Values in parentheses show difference from LLaVA 1.5 baseline.
@@
Both methods consistently outperform the original LLaVA 1.5 model, particularly as spatial complexity (the number of objects) increases (Figure 5), validating our hypothesis about the critical roles of embedding normalization and local spatial detail preservation.
6.3 Results on Standard Benchmarks
@@
To assess generalization, we evaluate our methods on established benchmarks (Table 3) that are used in recent works [23, 32, 41, 24].
@@
7 Analysis
To deepen our interpretability-driven understanding, we analyzed how the combination of embedding normalization and earlier-layer feature extraction influences the attention mechanisms of VLMs.
7.1 Attention Visualization
We visualized the self-attention patterns from LLM to interpret the effects of each modification on spatial reasoning. Figure 7 provides representative attention maps across four illustrative queries. Each query compares three scenarios: the baseline (Original LLaVA 1.5), normalized embeddings, and normalized embeddings combined with multilayer features. We visualized the first text token of the target object in the answer, which are bold and italic words in Figure 7. More examples are in Appendix C.

### v2

@@ 5.4 A 2D Synthetic Spatial Benchmark (2DS)
Figure 5: 2DS examples. Left: two-object layouts; right: three/four-object layouts. Semantics are simple to focus on spatial relations.
Standard benchmarks often admit semantic shortcuts (object co-occurrence, language priors), masking spatial failure modes (Section 3 and Fig. 2). Our dataset 2DS removes these shortcuts: scenes contain colored shapes placed at controlled locations; queries require absolute (e.g., “bottom”) or relative (e.g., “left of”) relations, with semantics (color/shape) serving only as identifiers. We generate multiple meta-categories by number of objects (2–6), randomize positions within each, and enumerate questions across {color, shape, color+shape} $\times$ {absolute, relative}. Full construction details are in Appendix B. We use the same methods from Section 3 to test 2DS’s position effectiveness, and the results are reported in Appendix B.2. It shows a much higher sensitivity over position compared to other datasets.
6 Probing Spatial Awareness: A Controlled Study
Guided by the diagnostics above, we test two concrete hypotheses about why VLMs underuse spatial cues under perturbations:
H1 (Norm-induced position embedding weakness). A modality norm mismatch (vision $\gg$ text) inflates vision token norms and attenuates RoPE effects.
H2 (Mid-layer spatial richness). Spatial detail is concentrated in intermediate vision-encoder layers, and final layers are more semantic, less spatial (Wang et al., 2023; Jiang et al., 2024).
6.1 Interventions and Setup
We use LLaVA-1.5 (7B) (Liu et al., 2023) as the baseline, reproducing its two-stage training and data. We propose and evaluate the following two minimal interventions. From the derivative in equation 4, H1 predicts that reducing the vision norm can improve positional sensitivity. H2 predicts intermediate layers can add additional gains from richer spatial evidence.
+ Normalize (H1). RMS normalization (Zhang & Sennrich, 2019a) of vision embeddings after the linear projector and before feeding the LLM, calibrated to typical text-token norms (empirically: mean $\approx 0.83$, max $\approx 1.22$).
+ Normalize + Multilayer (H2). Concatenate intermediate vision features to the vision token dimension (layers 12, 16, 20, 24) following (Jiang et al., 2024), then apply the same normalization. This aims to inject fine-grained spatial signals while controlling scale. Multilayer adds a small compute overhead in theory with extra embedding expanded from 1024 to 4096 dimension. We showed empirically it is negligible in Appendix D.
6.2 Position Sensitivity Index
Table 2: Position Sensitivity Index (%).
Data LLaVA 1.5 + Normalize + Normalize + Multilayer
VQAv2 1.09 2.03 2.08
POPE 0.34 0.35 0.46
GQA 4.47 6.10 5.64
CV-Bench 2D 0.57 11.31 9.72
2DS 41.07 61.21 54.21
PSI could measure a two-fold problem – model and dataset. Table 2 compares both for PSI. If we fix one dataset, we can compare across the models, and vice versa. Given a baseline model, PSI isolates model’s order use: if shuffling vision tokens barely moves accuracy compared to the baseline, the model likely underutilizes spatial arrangement. We can also compare it across benchmarks to see which one is more sensitive to the position changes. Particularly, we find 2DS has a significantly higher PSI compared to other datasets across different models, which means that 2DS can test position embedding more effectively than existing benchmarks.
(1) Normalization substantially increases PSI across diverse benchmarks, consistent with our hypothesis H1: reducing norm skew restores RoPE leverage and raises order-use. (2) Multilayer occasionally trails normalization on PSI (e.g., CV-Bench 2D, 2DS), yet as shown in Table 3 and Appendix E it can yield higher accuracy, indicating that models can exploit spatial content from vision encoder features even if their explicit sensitivity to token permutations is moderate. In short, PSI probes order use, while accuracy reflects spatial competence.
Table 3: 2DS accuracy (%) by attribute/relation. $\uparrow$ higher is better. $\Delta$ vs. baseline in parentheses.
@@ 6.2 Position Sensitivity Index
6.2.1 Which attributes are important?
Table 3 showed that +Normalize chiefly helps categories where color anchors spatial relations (Color_abs./rel., Shape_color_*), suggesting that final-layer features retain strong color cues but weaker geometry. By contrast, +Normalize+Multilayer substantially lifts relative categories across the board, aligning with H2: intermediate features furnish finer geometry and local layout that compose with color cues. The largest jumps appear where both feature identity and geometry must be integrated.
On 2DS, +Normalize attains the highest PSI (61.21%) while +Normalize+Multilayer yields the highest accuracy (64.8%). This suggests that restoring positional leverage (H1) and supplying spatial content (H2) are complementary spatial information encoding. Their relative contributions differ by metric.
6.3 Cross Modality Balance: Where do attention heads look?
System prompt confound.
Prior work reports very low vision attention vs. text (Chen et al., 2024b; 2025). Our toolkit separates system prompt from user text and vision, revealing why. Text can break into two categories, the system prompt, and the user input. We compare those two and the vision attentions. As in Fig. 6, the static system prompt absorbs a large, image-invariant portion of attention (72.1%), masking vision use. It might act as an attention sink (Xiao et al., 2024). It provides a new explanation on why prior work reports very low vision attentions. When we exclude the system prompt from our analysis, a different picture emerges: in the original VLMs, the attention mass on vision tokens is significant higher than the user input text. However, we believe the apparent visual dominance is not a sign of a healthy mechanism. Instead, it is a ”brute force” consequence of the large vision token norms, as described in Section 4. The normalization can balance the vision and user input text attention as we show in Appendix F.
Figure 6: LLaVA 1.5 7B attention share average over COCO Validation dataset. Green is system prompt, orange is vision token, and blue is user input text. Left graph includes system prompt. Right graph just compares vision and input text.
Modality-integration patterns. To understand how the VLMs integrate the vision and text modality, we visualize the CMB with heatmaps across the layers and attention heads. It can give us a granular sense of what each attention head focuses on, vision or user input text. Figure 7 showed the CMB heatmap visualization average across 5k COCO val results.
With prompt excised for clarity, LLaVA-1.5 shows many vision-specialized heads scattered across depth. It aligns with the analysis in Section 4 that the vision tokens could dominate the attention calculation. +Normalize exhibits a vision to text progression: early layers lean on vision, deeper layers shift toward text, indicating more balanced cross-modal exchange once norm disparity is removed. +Normalize+Multilayer displays a striking stage-like pattern: the first two layers emphasize vision, an early-block pivot concentrates on text, and deeper layers re-amplify vision with smooth transition. We view it as a pipeline that extract, align, then integrate. We also provide more individual image visualization and more dettails of calculation in Appendix G.
[…]
The performance of models over standard benchmarks in Appendix E show that +Normalize+Multilayer is the best, then +Normalize, then original. It suggest original model’s “visual dominance” does not mean a better visual performance. With large vision activations in the residual stream (Sec. 4), the normalized state is pre‑aligned to the vision subspace, so queries favor vision keys by default, and high CMB reflects a directional bias towards vision, not use question conditioned retrieval. As our attention visualizations and entropy measures indicate (Appx. H), baseline vision maps are diffuse and often focus on the wrong target objects asked in text. In contrast, +Normalize+Multilayer’s staged vision-text-vision has the most focused (lowest‑entropy) attention maps and the best spatial accuracy.
Figure 8: RoPE sensitivity (higher = larger change under phase shift) across layers.
6.4 RoPE sensitivity: How does normalization help?
One key question we want to understand is whether the normalization helped or not. We already see that the normalization can effectively balance the hidden states in Section 4 Figure 4. However, to have a more indepth understanding, we want to know how RoPE directly changes vision. Our RoPE probe perturbs positional phase for vision keys only and measures (Fig. 8): (i) attention-level change $\Delta\alpha$, and (ii) the normalized logit change $\Delta g_{V}$, which factors out attention imbalance as we observed from CMB results. We show the RoPE probe with $\Delta=1$ applied to the vision keys, and we report the RoPE sensitivity measures over COCO dataset average in Figure 8. We see both $\Delta\alpha_{V}$ and $\Delta g_{V}$ increase the sensitivity in early 0-5 layers over RoPE changes on vision tokens.
Both +Normalize and +Normalize+Multilayer show markedly higher sensitivity than baseline in early layers (0–5). (Fig. 8). This matches H1: reducing norm skew raises sensitivity of the RoPE effects. We believe it shows that normalization helps restoring vision token positional information from RoPE, especially in early layers. Position embedding increase the spatial channel information, improving both order-use (PSI) and task competence (2DS).
7 Discussion and Future Work
What does this reveal about VLM internal mechanisms? Spatial signals enter through (i) vision-encoder features and (ii) the LLM’s positional mechanism (RoPE). Our probes indicate that both are necessary but they play different roles. Scale matching (Sec. 4) restores order sensitivity (higher PSI and RoPE probes). Exposing intermediate vision layers supplies finer geometry and local layout that raise spatial competence even when PSI changes modestly. The division of labor therefore shifts with design: +Normalize leans on RoPE-driven order, and let the LLM do more works on spatial information. +Normalize+Multilayer delegates more spatial work to the vision encoder.
Why are raw vision attentions “low”? Layer-averaged attention under-counts visual use because the static system prompt behaves like an attention sink. Once we excluded it for analysis only, many attention heads are vision-preferring (Section 6.3), especially early/mid layers. Normalization rebalanced this imbalance caused by residual norm (vision share drops from $\sim$75% to $\sim$54% as shown in Appendix F). Results suggest a more balanced vision and text could perform better, possibly due to a better question conditioned retrieval. Vision attention are also less diffused and focus less on the wrong object as shown in Appendix H.
[…]
Limitations and Implications Our normalization is blunt and can reduce dynamic range, especially for 3D cues. It reduces the range of the expression and might lead to reduced 3D performance. Future work should consider a more sophisticated normalization alignment.
Our work is limited to the LLaVA style model, which has a pre-trained vision encoder align with a pre-trained LLM through a projection layer. Recent VLMs train vision encoders from scratch and might not have the problem we mentioned (Bai et al., 2025; Grattafiori et al., 2024).
8 Conclusion
We show how VLMs use spatial information, and we provide a compact interpretability toolkit to make this visible and measurable. Best to our knowledge, we are the first to connect embedding–norm skew to diminished positional sensitivity in the LLM (RoPE), with analysis and experiments supporting the link. This mechanistic understanding reveals how modality-specific scales can inadvertently suppress cross-modal information flow. We introduce four simple, targeted tools for spatial information measurements: PSI for order use, CMB for head-level modality balance with prompt isolation, a RoPE sensitivity probe for intrinsic phase response, and 2DS, a controlled spatial benchmark. We showed two minimal changes and clarified design levers: Normalize (scale-match vision to text) restores RoPE leverage and raises PSI and RoPE sensitivity; Multilayer adds mid-layer geometry and boosts performance accuracy. We believe those tools and insights are valuable for future VLMs design considerations.
Acknowledgments
The work is supported by the National Science Foundation (NSF) through Awards #2131186 (CISE-MSI), #1827505 (PFI), and the US Air Force Office of Scientific Research (AFOSR) via Award #FA9550-21-1-0082. The work is also supported by a College-wide Research Vision (CRV) Fund from the CCNY Provost’s Office and the Google CyberNYC Initiative. This work used Google Cloud through the CloudBank project, which is supported by National Science Foundation Award #1925001. We specifically thank Dr. Ping Ji’s generous support for computing resources.
Appendix A Theoretical Analysis of Large-Norm Effects on RoPE
We consider a pre-norm Transformer block (as used in LLaVA-style decoders). For token state $h^{(l)}\!\in\!\mathbb{R}^{d}$ at layer $l$,
$\underbrace{a^{(l)}(\phi)}_{\text{attn out}}\;=\;\mathrm{Attn}\!\big(\mathrm{RMSNorm}(h^{(l)}),\phi\big),\qquad h^{(l+1)}\;=\;\underbrace{r^{(l)}}_{\text{residual carry}}\;+\;a^{(l)}(\phi),$
where $\phi$ denotes the RoPE phase and $r^{(l)}\!=\!h^{(l)}$ (the pre-attention input).
Keys and queries are formed from per-token RMS-normalized inputs; RoPE applies a norm-preserving rotation to $q,k$ (Su et al., 2023). However, the residual is added to every normalization results even after the post layer norm. It causes the unbalanced vision token norm to keep passing on.
Residual-scale suppression (directional lemma).
With $h^{(l+1)}=r^{(l)}+a^{(l)}(\phi)$ and $u(\phi)=h^{(l+1)}/\|h^{(l+1)}\|$. With some quotient rule and replacing $\frac{\partial\|h^{(l+1)}\|}{\partial\phi}=\frac{(h^{(l+1)})^{T}}{\|h^{(l+1)}\|}\,\frac{\partial h^{(l+1)}}{\partial\phi}$ we can have
$\frac{\partial u}{\partial\phi}=\frac{(I-uu^{\top})}{\|h^{(l+1)}\|}\,\frac{\partial a^{(l)}}{\partial\phi}.$
When $\|r^{(l)}\|\!\gg\!\|a^{(l)}\|$ the magnitude of $\partial u/\partial\phi$ is suppressed by $\sim 1/\|r^{(l)}\|$ in the vision token part. Particularly, the current layer’s sensitivity is proportional to the attention sensitivity of the previous layer divided by a residual norm. A large vision token norm would mean much less effective position sensitivity. One of the key terms is the attention level sensitivity. It complements attention-level analyses by explaining why even nontrivial attention shifts may yield small representational changes when the residual dominates.
B. Attention-level phase sensitivity (group form)
Within a given head, let logits be $\ell_{j}=\langle q^{\prime},k^{\prime}_{j}\rangle/\sqrt{d}$, and define attention weights $\alpha_{j}=\mathrm{softmax}(\ell)_{j}=\frac{exp(\ell_{j})}{\sum_{i}exp(l_{i})}$. Partition keys into vision $V$ and text $T$, and define group masses
$\alpha_{V}=\sum_{v\in V}\alpha_{v},\qquad\alpha_{T}=1-\alpha_{V}.$
Let the group-average logit derivatives be
$g_{V}=\sum_{v\in V}\frac{\alpha_{v}}{\alpha_{V}}\frac{\partial\ell_{v}}{\partial\phi},\qquad g_{T}=\sum_{t\in T}\frac{\alpha_{t}}{\alpha_{T}}\frac{\partial\ell_{t}}{\partial\phi}.$
A standard calculation with some quotient rule from softmax yields the single vision key given a fixed query:
$\frac{\partial\alpha_{v}}{\partial\phi}=\alpha_{v}\left(\frac{\partial\ell_{v}}{\partial\phi}-\sum_{k}\alpha_{k}\frac{\partial\ell_{k}}{\partial\phi}\right)$ (5)
We can use the above notation and get the group-level sensitivity by aggregating all vision attentions then substitute with $\alpha_{V},\alpha_{T},g_{V},g_{T}$
$\frac{\partial\alpha_{V}}{\partial\phi}$ $=\alpha_{V}\left(g_{V}-\sum_{k}\alpha_{k}\frac{\partial\ell_{k}}{\partial\phi}\right)$ (6)
$=\alpha_{V}\left(g_{V}-(\sum_{v\in V}\alpha_{v}\frac{\partial\ell_{v}}{\partial\phi}+\sum_{t\in T}\alpha_{t}\frac{\partial\ell_{t}}{\partial\phi})\right)$ (7)
$=\alpha_{V}\left(g_{V}-(\alpha_{V}g_{V}+\alpha_{T}g_{T})\right)$ (8)
$=\alpha_{V}\alpha_{T}\left(g_{V}-g_{T}\right)$ (9)
Eq. equation LABEL:eq:group-derivative-appendix is an attention-level statement, and we can directly measure the change without the need to consider the previous layer.
Small-phase probe.
If we rotate only vision keys by a small phase $\Delta$ (queries fixed), then $g_{T}$ is unchanged and to first order
$\Delta\alpha_{V}\;\approx\;\alpha_{V}(1-\alpha_{V})\,\Delta g_{V},$ (11)
This factorization underlies our probe: $\Delta\alpha_{V}$ grows when (i) the balance factor $\alpha_{V}(1-\alpha_{V})$ is large (less saturation) and/or (ii) the intrinsic phase response $\Delta g_{V}$ is large.
Empirically we can capture $q,k$ inside attention and conduct position perturbation, so we can measure $\Delta\alpha_{V}$ and $\Delta g_{V}$ with attention-level sensitivity. The effective positional change is modulated by the residual. Reducing the interface scale (our +Normalize) decreases $\|r^{(\ell)}\|$ for vision tokens in early layers, mitigating residual-scale suppression and making attention-level phase shifts consequential for subsequent computation. It also confirmed with our observation that +Normalize increases early layers’ position sensitivity in Section 6.4.
Appendix B 2DS Dataset Details
Example Queries including:
  • “What color is at the bottom of the image?” (Color, absolute position)
  • “Is the circle below the square?” (Shape, relative position)
B.1 Dataset Construction Details
B.2 How effective can 2DS measure the position embedding sensitivity
As the preliminary experiments in Section 3 showed that most of the data retain the performance even with random permutation, we can use the same experiments to understand the 2DS results.
Permutation
Table 4 show the 2DS with permutation of the input tokens. We also report the difference which is used to compute the PSI in Table 2. We can see the +Normalized version is more susceptible to the token permutation. As the baseline, we also test with the Vicuna-7B to show the model performance is well above the random guess.
Table 4: For 2DS, we report accuracy on original (Orig.) and permuted (Perm.) token arrangements with difference (Diff.).
Model 2DS (%)
Orig. Perm. Diff.
Vicuna-7B 15.97
LLaVA 1.5 56.63 33.37 -23.26
+ Norm 59.30 23.00 -36.30
+ Norm + Multi 64.80 29.67 -35.13
Compression
Another method to test the 2DS is through compression, which eventually reduce position embeddings effectiveness. If the vision encoder can successfully carry over most of the spatial information, then the compression would not hurt performance significantly. However, if the position embedding is crucial, then there should be a significant decrease. We show the exact same plot but with 2DS through the compression in Figure 9.
Here we see a significantly more drop in the 2DS performance. We believe those two measures suggest that position embedding has an important role for the 2DS performance.
Appendix C Training Details
We follow the training recipe described in LLaVA 1.5 7B (Liu et al., 2023), consisting of two stages: pretraining and instruction tuning. All experiments utilize 8 Nvidia A100 GPUs. Learning rates for each training stage are detailed in 5, with all other hyperparameters and datasets identical to (Liu et al., 2023).
Table 5: Learning rate of training
Model Pre-train Finetune
LLaVA 1.5 7B 1e-3 2e-5
+Normalize 1e-3 2e-5
+Normalize+Multilayer 1e-3 2e-5
Appendix D Analysis of Computational Cost and Inference Time
In this section, we analyze the computational overhead of our two proposed methods to evaluate their practical efficiency.
Method 1: +Normalize. Our first method, which applies L2 normalization to the final-layer vision tokens, introduces zero additional parameters or floating-point operations (FLOPs) during inference. The normalization is a lightweight mathematical operation on the existing embedding vector and its computational cost is negligible.
Method 2: +Normalize+Multilayer. Our second method enriches the visual representation by concatenating features from the intermediate and final layers of the vision encoder. This modification affects the initial visual projection layer. In the original LLaVA, the projector is an MLP that maps final 1024-dimensional feature vector to 4096-dimensional LLM embedding space. While our proposed method concatenates four 1024-dimensional feature vectors and inputs them to the projector, producing 4096-dimensional vector. While this moderately increases the number of parameters in this single projection layer, this component represents a very small fraction of the total parameters in the multi-billion parameter VLM.
Empirical Inference Time. To assess the real-world impact on performance, we benchmarked the inference speed of the baseline model against our two proposed methods. The experiment was conducted on a single NVIDIA RTX 4090 GPU, measuring the average time to process 500 samples from the GQA dataset validation split. All three models are within the same inference time $(0.137s\pm 0.002s)/sample$, empirically there are trivial compute cost differences.
Appendix E Benchmark Performance
E.1 2DS Performance
(a) Accuracy comparison across varying numbers of objects. Our interpretability-informed adjustments yield consistent improvements, especially as spatial complexity increases.
(b) Accuracy comparison across different categories of 2DS dataset. We see consistent improvements across categories.
Figure 10: Model performance across increase of objects and different sub-categories
Table  3 summarizes accuracy across distinct spatial query categories, and Figure  10(b) illustrates performance across varying numbers of objects, clearly highlighting the effectiveness of our interventions.
@@ Appendix E Benchmark Performance
Both methods consistently outperform the original LLaVA 1.5 model, particularly as spatial complexity (the number of objects) increases (Figure  10(a)), validating our hypothesis about the critical roles of embedding normalization and local spatial detail preservation.
E.2 Results on Standard Benchmarks
@@
Table 7: Performance comparison on CV-Bench 2D benchmarks. We report accuracy on original (Orig.) and permuted (Perm.) token arrangements and PSI for each of the three tasks: overall 2D spatial reasoning, counting, and relation. We also report evaluation using Vicuna-7B on top without image input but just text input as the random baseline.
Model CV-Bench 2D (%)
Overall Count Relation
Orig. Perm. PSI Orig. Perm. PSI Orig. Perm. PSI
Vicuna-7B 31.36 12.31 54.46
LLaVA 1.5 56.59 56.26 0.58 52.07 52.37 -0.58 62.05 60.97 1.74
+ Norm 59.91 53.13 11.32 52.81 51.14 3.16 68.51 55.54 18.93
+ Norm + Multi 58.69 52.99 9.71 52.08 49.49 4.97 66.72 57.23 14.22
To assess generalization, we evaluate our methods on established benchmarks (Table  6) that are used in recent works (Liu et al., 2023; Tong et al., 2024a; Zhu et al., 2024; Liu et al., 2024).
@@
Specifically, we investigate the subcategories of the CV-Bench2D as shown in Table 7. We are curious why the permutation drop was not significant. As shown in Table 7, we can check the PSI index to see whether the model is sensitive to that task, and Count is not really a spatial task since there are marginal effects when changing spatial information. In comparison, we see the Relation task is more spatial, and the +Normalize has a much better performance in that category.
Appendix F System Prompt Attention
We visualize the attention portion when inference with and without system prompt for all 3 models in Figure 11. We observe that the system prompt absorbs the most attention shares. Even when we exclude the system prompt during the calculation, it might skew the probability mass and changes the distribution of a much smaller vision and user input text attentions. Therefore we also demonstrate vision and user input text attention inferenced without the system prompt on the right column of the Figure 11.
Figure 11: Visualization of system prompt, vision, and user input text attention shares. Left 2 columns are inferenced with the system prompt, then calculate the portion of the vision and user input text. Right column is inferenced without the system prompt. We see a much more balanced vision and user input text attention share in the normalized and multilayer version.
Appendix G Attention Head Visualization
Figure 12: Visualization of each image’s attention head over layers. We inference without the system prompt. CMB directly calculated from vision and text attention.
Figure 13: Visualization of each image’s attention head over layers. We inference with the system prompt. CMB take values from vision and text attentions and normalize them with the sum of two.
Appendix H Attention Visualization
We visualized the self-attention patterns from LLM to interpret the effects of each modification on spatial reasoning. Figure  14 provides representative attention maps across four illustrative queries. Each query compares three scenarios: the baseline (Original LLaVA 1.5), normalized embeddings, and normalized embeddings combined with multilayer features. We visualized the first text token of the target object in the answer, which are bold and italic words in Figure  14. More examples are in H.2.